In a cycle of copper experiment, a student first reacts a piece of copper metal with nitric acid to produce copper(II) nitrate (Cu(NO)2)

(Cu(NO3)2) solution. The student then performs various reactions which transform the copper ions into a series of different copper compounds and complexes. Finally, the last reaction reduces the copper ions back to elemental copper metal. Copper atoms are conserved throughout the process.

If the initial step of the experiment produces 8.20 mL

8.20 mL of 1.21 M Cu(NO3)2,

1.21 M Cu(NO3)2, what is the theoretical yield of solid copper (Cu)

(Cu) that can be recovered at the end of the experiment?


1
Expert's answer
2021-04-07T11:32:54-0400

"Moles = \frac{1.21}{8.20×10^{-3}}" =147.56 moles



Molar mass: 187.56 g/mol


Mass = 147.56 ×187.56 = 27676.35grams